# SimTxtSeg: Weakly-Supervised Medical Image Segmentation with Simple Text Cues

### v1

@@ Abstract
Weakly-supervised medical image segmentation is a challenging task that aims to reduce the annotation cost while keep the segmentation performance. In this paper, we present a novel framework, SimTxtSeg, that leverages simple text cues to generate high-quality pseudo-labels and study the cross-modal fusion in training segmentation models, simultaneously. Our contribution consists of two key components: an effective Textual-to-Visual Cue Converter that produces visual prompts from text prompts on medical images, and a text-guided segmentation model with Text-Vision Hybrid Attention that fuses text and image features. We evaluate our framework on two medical image segmentation tasks: colonic polyp segmentation and MRI brain tumor segmentation, and achieve consistent state-of-the-art performance.
@@ 1 Introduction
The Segment Anything Model (SAM)[12], a general visual foundation segmentation model, has garnered widespread attention due to its remarkable segmentation and robust zero-shot generalization capabilities. Although SAM has been trained using large-scale data with pixel-level labels, its performance for medical image segmentation is unsatisfactory due to the lack of reliable clinical training data. Consequently, many researchers have fine-tuned SAM specifically in medical domains[17, 31, 13, 4], including full fine-tuning and parameter-efficient tuning, achieving promising performance. Nonetheless, models based on SAM still require providing manual visual prompts (e.g. point and box prompts) for each image, increasing the difficulty and time required for expert physician to make annotations. Therefore, we aim to explore a novel and automatic approach only using simple text cues to accomplish weakly-supervised medical image segmentation, through equipping SAM with language-to-vision prompt converter. Moreover, to further conduct cross-modal fusion, we seek to better integrate text cues into the target visual segmentation model, ultimately enhancing the effect of language-driven segmentation performance.
@@ 2.2 Textual-to-Visual Cue Converter
Once the Textual-to-Visual Cue Converter is pre-trained, we can straightforwardly transfer the text prompts into visual prompts for any new dataset within the same medical domain. Then we employ SAM as our pseudo-masks generator with the visual prompts $B$, configuring the confidence threshold at 0.25. In our study, both the vanilla SAM and SAM-med2d are experimented.
@@ 2.3 Text-Guided Segmentation with Text-Vision Hybrid Attention
Vision Encoder: Given an image $I$, we choose ConvNext-Tiny[28] as our vision encoder $\Phi_{vision}$: $v_{i}=\Phi_{vision}(I),$ where $i$ refers to the $i$-th layer in the backbone. We extract its first four layers of output for feature fusion, which are defined as $v_{1}\in\mathbb{R}^{H/4\times W/4\times C1}$, $v_{2}\in\mathbb{R}^{H/8\times W/8\times C2}$, $v_{3}\in\mathbb{R}^{H/16\times W/16\times C3}$, $v_{4}\in\mathbb{R}^{H/32\times W/32\times C4}$.
Text Encoder: Given a sentence $T$, we take BERT-BASE[5] as our tokenizer and text backbone, and take its last embedding $t\in\mathbb{R}^{l\times C}$, where $l$ is the length of token, and $C$ refers to the feature dimension. $t=\Phi_{text}(tokenize(T)).$
@@
Let $v^{h}$ represents the high-level feature from the previous decoder layer ($v^{h}=v_{4}$ in the first decoder), and $v^{l}$ represents the low-level feature from the corresponding encoder layer, we upsample $v^{h}$ and concatenate it with $v^{l}$, then obtain the output $f_{v}$ as:
$f_{v}=Concat(Upsample(v^{h}),v^{l}).$ (3)
@@
where $\operatorname{MLP}$ is a multiple layer perception, containing a 1 $\times$ 1 convolution layer, a GELU activation function and a linear layer. $f_{t}\in\mathbb{R}^{L_{j}\times C_{j}}$ is the output text embedding, $L_{j},C_{j}$ represent the length and channel number of the output token in the $j$-th decoder layer. Also, the image embedding’s shape is projected into $HW\times C_{j}$, consistent with $f_{t}$.
@@
where $MHSA(\cdot)$ refers to Multi-Head Self Attention. Second, a text-to-vision attention is applied, which means cross-attention from text (text position embedding as query) to the image embbeding (image position as key, $f_{v}$ as value). The text-to-vision cross-attention map can be formulated as:
$\mathcal{M}_{tv}\left(f_{t},f_{v}\right)=\sigma\left(\frac{Pos_{t}(f_{t})W_{q}\left[Pos_{v}(f_{v}^{\prime})W_{k}\right]^{T}}{\sqrt{d_{m}}}\right)\left(f_{v}^{\prime}W_{v}\right)$ (6)
where $\sigma$ is the Softmax function, $Pos_{v},Pos_{t}$ refer to the position encoder of text and image embedding, $W_{q},W_{k},W_{v}$ are the learnable weight matrices used to project $f_{v}^{\prime}$ and $f_{t}$ to different feature subspaces. $d_{m}$ is the dimension of query and key. Then, a norm-and-add layer is applied. The text-to-vision cross-attention process is shown in Eq. 7:
@@ 3.1 Experiment Setup
MRI Brain Tumor Dataset: For brain tumor segmentation, we utilize LGG Segmentation Dataset[3] from The Cancer Imaging Archive, which comprises 3,929 brain MRI images with a uniform size of $256\times 256$ pixels. Other settings remain consistent with the Polyp datasets.
@@
Implementation Details: To pre-train the Textual-to-Visual Cue Converter, we employed Adam optimizer with an initial learning rate of $2\times 10^{-4}$, weight decay of $1\times 10^{-4}$, a batch size of 4, and trained it for 100 epochs. As for the parameter scheduler, we adopted both LinearLR and MultiStepLR. To train the text-guided segmentation model with Text-Vision Hybrid Attention, we freeze the text branch parameters and employ ConvNeXt as the vision backbone, with an input image size of 384. The learning rate adjustment strategy is ReduceLROnPlateau. All the methods are implemented using PyTorch, accelerated by an NVIDIA 4090 Ti GPU.
@@ 3.2 Comparisons with the State-of-the-Art Methods
Comparison results against seven state-of-the-art methods are reported in Table 1. These methods fall into two categories: three fully-supervised models (ResUNet[6], PraNet[7], and Ariadne’s Thread[32]) and four weakly-supervised models with different label levels (WeakPolyp[26], BoxPolyp[27], Boxshrink[9], and S2ME[22]). We compared the segmentation performance of these SOTAs with our proposed pseudo-label generator(Pseudo-L:TVCC+SAM) and the final weakly-supervised model(SimTxtSeg-w-TVHA). It is observed that the generated pseudo-label quality is roughly on par with that of the SOTA fully-supervised models, with even a slight edge on the polyp dataset, and our final segmentation performance surpasses other SOTA weakly-supervised models. Among different kinds of weak supervision cue, our text-based cue is the weakest annotation without any spatial labeling and has the lowest cost compared to visual cues like boxes and scribbles which still cost a lot. Specifically, on the polyp dataset, we achieve a +1.38% improvement in mDice and a +3.36% improvement in mIoU. Moreover, on the brain tumor dataset, our method achieves a +4.1% improvement in mDice and a +3.94% improvement in mIoU. Qualitative comparison of segmentation performance is visualized in Fig. 3.
@@ 3.3 Ablation Study
Impact of prompt types. We evaluated class name and sentence as text cues during training textual-to-visual cue converter and compared their effectiveness for pseudo-mask generation on SAM-huge. Also, we tested the performance by the original GroundingDINO[15] with class name prompt. As Table 2 shows, training the textual-to-visual cue converter with sentences (e.g. A polyp is an anomalous oval-shaped small bump-like structure.) tends to yield slightly better results than training it with class names (e.g. polyp), since the converter generates better pseudo boxes, achieving 0.8010 mAP for polyp and 0.7480 mAP for brain tumor. The original GroundingDINO fails to generate useful pseudo masks.

### v3

@@ Abstract
Weakly-supervised medical image segmentation is a challenging task that aims to reduce the annotation cost while keep the segmentation performance. In this paper, we present a novel framework, SimTxtSeg, that leverages simple text cues to generate high-quality pseudo-labels and study the cross-modal fusion in training segmentation models, simultaneously. Our contribution consists of two key components: an effective Textual-to-Visual Cue Converter that produces visual prompts from text prompts on medical images, and a text-guided segmentation model with Text-Vision Hybrid Attention that fuses text and image features. We evaluate our framework on two medical image segmentation tasks: colonic polyp segmentation and MRI brain tumor segmentation, and achieve consistent state-of-the-art performance. Source code is available at: https://github.com/xyx1024/SimTxtSeg.
@@ 1 Introduction
The Segment Anything Model (SAM)[12], a general visual foundation segmentation model, has garnered widespread attention due to its remarkable segmentation and robust zero-shot generalization capabilities. Although SAM has been trained using large-scale data with pixel-level labels, its performance for medical image segmentation is unsatisfactory due to the lack of reliable clinical training data. Consequently, many researchers have fine-tuned SAM specifically in medical domains[17, 31, 13, 4], including full fine-tuning and parameter-efficient tuning, achieving promising performance. Nonetheless, models based on SAM still require providing manual visual prompts (e.g. point and box prompts) for each image, increasing the difficulty and time required for expert physicians to make annotations. Therefore, we aim to explore a novel and automatic approach only using simple text cues to accomplish weakly-supervised medical image segmentation, through equipping SAM with a language-to-vision prompt converter. Moreover, to further conduct cross-modal fusion, we seek to better integrate text cues into the target visual segmentation model, ultimately enhancing the effect of language-driven segmentation performance.
@@ 2.2 Textual-to-Visual Cue Converter
Once the Textual-to-Visual Cue Converter is pre-trained, we can straightforwardly transfer the textual prompts into visual prompts for any new dataset within the same medical domain. Then we employ SAM as our pseudo-masks generator with the visual prompts $B$, configuring the confidence threshold at 0.25. In our study, both the vanilla SAM and SAM-med2d are experimented.
@@ 2.3 Text-Guided Segmentation with Text-Vision Hybrid Attention
Vision Encoder & Text Encoder: Given an image $I$, we choose ConvNext-Tiny[28] as our vision encoder $\Phi_{vision}$: $v_{i}=\Phi_{vision}(I),$ where $i$ refers to the $i$-th layer in the backbone. We extract its first four layers of output for feature fusion, which are defined as $v_{1}\in\mathbb{R}^{H/4\times W/4\times C1}$, $v_{2}\in\mathbb{R}^{H/8\times W/8\times C2}$, $v_{3}\in\mathbb{R}^{H/16\times W/16\times C3}$, $v_{4}\in\mathbb{R}^{H/32\times W/32\times C4}$. Given a sentence $T$, we take BERT-BASE[5] as our tokenizer and text backbone, and take its last embedding $t\in\mathbb{R}^{l\times C}$, where $l$ is the length of token, and $C$ refers to the feature dimension. $t=\Phi_{text}(tokenize(T)).$
@@
Let $v^{h}$ represents the high-level feature from the previous decoder layer ($v^{h}=v_{4}$ in the first decoder), and $v^{l}$ represents the low-level feature from the corresponding encoder layer, we upsample $v^{h}$ and concatenate it with $v^{l}$, then obtain the output $f_{v}$.
@@
where $\operatorname{MLP}$ contains a 1 $\times$ 1 convolution layer, a GELU activation function and a linear layer. $f_{t}\in\mathbb{R}^{L_{j}\times C_{j}}$ is the output text embedding, $L_{j},C_{j}$ represent the length and channel number of the output token in the $j$-th decoder layer. Also, the image embedding’s shape is projected into $HW\times C_{j}$, consistent with $f_{t}$.
@@
where $MHSA(\cdot)$ refers to Multi-Head Self Attention. Second, a text-to-vision attention is applied, which means cross-attention from text (text position embedding as query) to the image embbeding (image position as key, $f_{v}$ as value). Then, a norm-and-add layer is applied. The text-to-vision cross-attention process is shown in Eq. 5:
@@ 3.1 Experiment Setup
MRI Brain Tumor Dataset: For brain tumor segmentation, we utilize LGG Segmentation Dataset[3] from The Cancer Imaging Archive, which comprises 3,929 brain MRI images. Other settings remain consistent with the Polyp datasets.
@@
Implementation Details: To pre-train the Textual-to-Visual Cue Converter, we employed Adam optimizer with an initial learning rate of $2\times 10^{-4}$, weight decay of $1\times 10^{-4}$, a batch size of 8, and trained it for 100 epochs. As for the parameter scheduler, we adopted both LinearLR and MultiStepLR. To train the text-guided segmentation model with Text-Vision Hybrid Attention, we freeze the text branch parameters and employ ConvNeXt as the vision backbone, with an input image size of 384. The learning rate adjustment strategy is ReduceLROnPlateau. All the methods are implemented using PyTorch, accelerated by an NVIDIA 4090 Ti GPU.
@@ 3.2 Comparisons with the State-of-the-Art Methods
Comparison results against seven state-of-the-art methods are reported in Table 1. These methods fall into two categories: three fully-supervised models (ResUNet[6], PraNet[7], and Ariadne’s Thread[32]) and four weakly-supervised models with different label levels (WeakPolyp[26], BoxPolyp[27], Boxshrink[9], and S2ME[22]). We compared the segmentation performance of these SOTAs with our proposed pseudo-label generator(Pseudo-L:TVCC+SAM) and the final weakly-supervised model(SimTxtSeg-w-TVHA). It is observed that the generated pseudo-label quality is roughly on par with that of the SOTA fully-supervised models, with even a slight edge on the polyp dataset, and our final segmentation performance surpasses other SOTA weakly-supervised models. Among different kinds of weak supervision cue, our text-based cue is the weakest annotation without any spatial labeling and has the lowest cost compared to visual cues like boxes and scribbles which still cost a lot. On the polyp dataset, we achieve a +1.38% improvement in mDice and a +3.36% improvement in mIoU. Moreover, on the brain tumor dataset, our method achieves a +4.1% improvement in mDice and a +3.94% improvement in mIoU. Qualitative comparison of segmentation performance is visualized in Fig. 3.
@@ 3.3 Ablation Study
Impact of prompt types. We evaluated class name and sentence as text cues during training textual-to-visual cue converter and compared their effectiveness for pseudo-mask generation on SAM-huge. Also, we tested the performance by the original GroundingDINO[15] with class name prompt. As Table 2 shows, training the textual-to-visual cue converter with sentences (e.g. A polyp is an anomalous oval-shaped small bump-like structure.) tends to yield slightly better results than training it with class names (e.g. polyp), since the converter generates better box cues, achieving 80.1% mAP for polyp and 74.8% mAP for brain tumor. The original GroundingDINO fails to generate useful pseudo masks.
@@ 4.0.1 Acknowledgements
[…]
The authors have no competing interests to declare that are relevant to the content of this article.